The number of road construction projects that take place at any one time in a certain city follows a Poisson distribution with a mean of 7. Find the probability that more than four road construction projects are currently taking place in the city


1
Expert's answer
2022-04-26T14:08:37-0400

Let "X=" the number of road construction projects: "X\sim Po(\lambda)."  

Given "\lambda=7."


"P(X>4)=1-P(X\le 4)"

"=1-P(X=0)-P(X=1)"

"-P(X=2)-P(X=3)-P(X=4)"

"=1-\dfrac{e^{-\lambda}\cdot\lambda^0}{0!}-\dfrac{e^{-\lambda}\cdot \lambda^1}{1!}"

"-\dfrac{e^{-\lambda}\cdot\lambda^2}{2!}-\dfrac{e^{-\lambda}\cdot \lambda^3}{3!}-\dfrac{e^{-\lambda}\cdot \lambda^4}{4!}"

"=1-e^\n{-7}(1+7+49\/2+343\/6+2401\/24)"

"\approx 0.82701"